OpenStudy (anonymous):

EASY MEDAL 5. What is the approximate hydronium ion concentration and hydroxide ion concentration in a cup of tea? Which is higher? (Refer back to the chart of pH if you need to.)

OpenStudy (anonymous):

i could use some help here...

OpenStudy (anonymous):

do u know the pH of tea?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

and whts that

OpenStudy (anonymous):

4-6

OpenStudy (anonymous):

now to convert pH into [H+] u need to take anti loge of negative pH

OpenStudy (anonymous):

what do you mean "anti loge" ?

OpenStudy (anonymous):

*anti-log

OpenStudy (anonymous):

ok so how do i take anti log of negative pH

OpenStudy (anonymous):

\[[H^+] = 10^{-4}\] \[[H^+] =1 * 10^{-4}\] if u find [OH-] then pOH= 14 - pH pOH= 14 - 4 pOH= 10 so \[[OH^+] = 10^{-10}\] \[[OH^+] =1 * 10^{-10}\] that means hydronium ion concentration in cup of tea is more

OpenStudy (anonymous):

got it??

OpenStudy (anonymous):

now u can also prove it for pH=6